OpenStudy (anonymous):

is √6 a rational or irrational ? any why

OpenStudy (anonymous):

irrational because it's not of the form a/b where a,b real numbers

OpenStudy (anonymous):

are you trying to prove it?

OpenStudy (saifoo.khan):

irrational, b.c. it dosnt have 2 real factors.

OpenStudy (anonymous):

let's assume it's rational... then: \[\sqrt{6}\] is positive => by Archimedian principle there is exist positive integers m & n such: \[\sqrt{6}=m/n\] or \[6n ^{2}=m ^{2}\] =>if assumption/ ration is true, then 6 is divisor of the left hand side (LHS) and also diviser right hand side (RHS) => 2 is also diviser of RHS & LHS this is only possible if 2 is a factor of m

OpenStudy (anonymous):

let m=2k (k= positive integer) => 4k^2 = 6 n^2 2k^2=3n^2 as 2 is a factor of LHS => 2 is also a factor of RHS. But this is only possible if 2 is a factor of n... but we assumed that \[\sqrt{6}\]=m/n presented in lowest term... if divided by 2 => not in lowest term => contradiction. =>\[\sqrt{6} \] is not a rational number